OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is y - x = 5?

OpenStudy (raden):

ax+by = c has the slope m = -a/b

OpenStudy (raden):

y - x = 5 here a = -1, b = 1, and c=5 so, m = -a/b = -(-1)/1 = 1/1 = 1

OpenStudy (anonymous):

Thanks

OpenStudy (raden):

welcome